OpenStudy (anonymous):

derivative of sqrt(1-x^2)

OpenStudy (anonymous):

\[\sqrt(1-x^2)\]

OpenStudy (saifoo.khan):

(1-2x)^1/2

OpenStudy (bahrom7893):

wrong

OpenStudy (bahrom7893):

-x/sqrt(1-x^2)

OpenStudy (anonymous):

can you explain your steps or write them out

OpenStudy (saifoo.khan):

Lol, i showed the first step.

OpenStudy (bahrom7893):

lol even your first step is wrong.. face it saif, you don't even know basic calc, so stop confusing people

OpenStudy (bahrom7893):

(1-x^2)^(1/2) (1/2)*(1-x^2)^(-1/2)*(-2x) -x(1-x^2)^(-1/2)=-x/(1-x^2)^(1/2)=-x/sqrt(1-x^2)

OpenStudy (anonymous):

you can also initially take y as the expression and square it and then do implicit diferrentiation

OpenStudy (anonymous):

somjit you above me :) ... Thanks bahrom